### v1

@@ II-C Importance
To mitigate these risks, systematic safety evaluation plays an essential role across the following key areas:
  • Risk Mitigation : Safety evaluations proactively identify and address vulnerabilities—such as the generation of toxic content or factual inaccuracies—before deployment. This process ensures that the model does not produce hate speech or biased outputs, safeguarding individuals and communities from harm while minimizing potential societal disruptions.
  • Ethical and Legal Compliance: Assessing the model for bias and adversarial vulnerabilities ensures that its behavior aligns with ethical principles and complies with relevant legal frameworks. This evaluation is particularly critical in high-stakes settings, such as healthcare, law, and education, where unethical or non-compliant outputs can lead to serious consequences.
  • User Trust Enhancement: A well-executed safety evaluation helps prevent the generation of misleading or inaccurate content, thereby fostering trust by ensuring users receive reliable and accurate information. Building trust is essential to encouraging the widespread adoption and acceptance of LLMs technologies across various sectors.
  • System Security Improvement: Evaluating for issues such as toxicity, bias, and truthfulness provides continuous feedback to developers, guiding iterative improvements throughout the model’s lifecycle. This feedback loop ensures the system remains secure, reliable, and aligned with evolving standards, facilitating long-term technical innovation and sustainable development.
This review seeks to fill existing research gaps by presenting a comprehensive and systematic framework for LLMs safety evaluation. Through the synthesis of current methodologies, tools, and metrics, this study aims to provide researchers with a clear overview of the state of the field, helping to reduce redundant efforts and accelerate progress. Moreover, we emphasize the need to promote standardization in safety evaluation practices, encouraging more researchers to engage with this critical area to advance the responsible and sustainable development of LLMs technologies.
@@ III WHAT TO EVALUATE
This section provides an overview of various perspectives for evaluating the safety of LLMs, including toxicity, robustness, ethics, bias and fairness, truthfulness, and other specific downstream tasks. It aims to illustrate what specific dimensions should be considered in evaluating and demonstrating the safety of LLMs, offering guidance for conducting comprehensive safety evaluations.
@@ III-A Toxicity
Toxicity refers to the presence of offensive, hateful, insulting, or harmful content, including incitement to violence, within the text generated by LLMs. Such toxic outputs not only pose risks of psychological harm to individual users but may also trigger broader social conflicts. Therefore, evaluating the potential of LLMs to generate toxic content has become a critical step in ensuring their safe deployment. In recent years, researchers have conducted extensive studies in this area:
ToxiGen: [2] highlights that existing toxicity detectors often over-rely on surface-level mentions of minority identities, leading to the neglect of subtle hate speech and the over-detection of benign expressions. To address this issue, the authors developed a large-scale dataset, ToxiGen, to accurately evaluate LLMs’ ability to generate adversarial and implicit toxic speech. To automatically generate these challenging texts, they introduced an adversarial classifier-in-the-loop decoding algorithm—ALICE. ALICE adjusts the toxicity of the generated text by comparing a toxicity classifier with the text generator during beam search decoding. By controlling machine generation in this manner, ToxiGen can encompass a broader range of implicit toxic texts than any previous human-written text resources, addressing a wider array of demographic groups. Experimental results demonstrate that several existing toxicity classifiers struggle to accurately distinguish between toxic and non-toxic outputs generated by ALICE. Moreover, comprehensive human evaluations show that ALICE-generated texts closely resemble human-produced content. This work offers a significant step forward in improving toxicity classifiers by providing both the dataset and the generation method.
@@
ToxiChat: [1] aims to address the challenge of detecting toxic content in real-world user-AI interactions. They note that most existing research on toxicity detection relies on benchmarks built from social media content, failing to capture the unique challenges of real-world conversations with AI systems. To fill this gap, the authors developed a new benchmark—ToxicChat, based on real user queries directed at open-source chatbots. To ensure data quality and reduce the burden of manual annotation, the authors employed an uncertainty-guided human-AI collaborative annotation approach. During the annotation process, they observed that chatbots are increasingly subjected to jailbreak attacks, underscoring the urgency of the current safety landscape. Furthermore, a systematic evaluation revealed that models trained on existing harmful content datasets perform poorly when applied to the ToxicChat domain, highlighting the need for enhanced safety measures in real-world user-AI interactions.
@@
LifeTox: [3] explores the challenge of detecting implicit harmful content within diverse life-advice scenarios. Implicit toxicity refers to a deeper, more concealed form of toxicity conveyed through linguistic features such as euphemisms, sarcasm, circumlocution, and metaphors, as well as extralinguistic knowledge like commonsense knowledge, world knowledge and social norms [100]. The authors argue that existing safety benchmark datasets are limited in diversity, primarily relying on red-team prompts, which often result in predictable and repetitive scenarios, failing to capture the complexity and variety of real-world situations.To address this gap, the authors introduced the LifeTox dataset, designed to identify implicit harmful content in various personal advice-seeking contexts, thereby improving the safety and reliability of the advice provided by LLMs. LifeTox was constructed by collecting posts from two Reddit forums focused on sharing life tips and seeking advice. Experimental results demonstrate that RoBERTa[101], fine-tuned on LifeTox, performs on par with or better than large language models in zero-shot harmful content classification, underscoring the effectiveness of LifeTox in tackling the complex challenges posed by implicit toxicity.
@@
PolygloToxicityPrompts: In contrast to previous toxicity assessments conducted in single-language contexts (such as English or Chinese), Jain et al. [5] explores the issue of LLMs generating harmful content in multilingual environments. They note that existing toxicity evaluation datasets predominantly focus on English or translations based on English benchmarks, leading to inadequate assessments of toxicity degradation in contemporary LLMs. To address this limitation, the authors propose PolygloToxicityPrompts, a large-scale multilingual benchmark containing 425K naturally occurring prompts across 17 languages. The prompts were generated by scraping documents from extensive corpora, with an average length of 400 GPT-4 tokens. During the evaluation, the authors employed the Perspective API[102] to assess the toxicity of the prompts and their generated content, calculating the average toxicity level of the models. Experimental results indicate significant differences in toxicity levels among LLMs across different languages, with an observed increase in toxicity as model size grows within the same family. Furthermore, models after instruction and preference-tuning exhibited lower toxicity compared to baseline models; However, the choice of preference-tuning method had a minimal impact on the models’ toxicity levels. The authors emphasize the need for further research on multilingual toxicity mitigation and the influence of model hyperparameters on toxicity.
@@
Red-Eval: [8] introduces a novel safety red teaming benchmark, RED-EVAL, to comprehensively evaluate the risks associated with harmful outputs generated by LLMs. RED-EVAL implements jailbreak attacks through a Chain of Utterances(CoU)-based prompt, facilitating a dialogue between two agents: a harmful agent Red-LM and a unsafe-helpful agent Base-LM. A harmful question is posed as the discourse for the Red-LM, and the model is instructed to generate a response according to the guidelines outlined in the prompt. Additionally, they propose the RED-INSTRUCT method, which enhances LLM safety through a two-phase approach: first, the CoU prompts are used to collect a dataset of harmful questions, termed HARMFULQA; Second, the model is safety-aligned by minimizing the negative log-likelihood of helpful responses while penalizing harmful ones. Experimental results show that RED-EVAL achieves a jailbreak success rate of 65% on GPT-4 and 73% on ChatGPT, highlighting its effectiveness in evaluating risks associated with harmful content generation.
In addition to the aforementioned studies, several other works have made significant contributions to the evaluation of LLMs toxicity. For instance, Cui et al. [4] collected questions and jailbreak templates from Bai et al. [103] and Liu et al. [104] to evaluate LLMs for utterance-level toxicity (referring to literally-toxic language with some typical words) and context-level toxicity (referring to that a harmless statement could be a toxic one when considered within its context.). All evaluated LLMs demonstrated an increase in toxicity from the utterance- to context-level evaluation. This performance gap may arise from literally non-toxic responses that inadvertently affirm toxic issues. Zou et al. [7] proposed greedy coordinate gradient-based search(GCG) to generate adversarial suffixes for malicious prompts, aiming to induce LLMs to produce harmful content. Additionally, the authors developed a new benchmark, AdvBench, designed for systematic evaluation based on harmful strings and harmful behaviors. Zhou et al.[6] conducted a comprehensive evaluation of ChatGPT using red teaming methods, highlighting that ChatGPT is vulnerable to prompt injection, leading to the generation of toxic content, and currently lacks the ability to detect toxicity at early stages.
@@ III-B Robustness
Robustness refers to the ability of LLMs to maintain stable and safe outputs when faced with noise (e.g., text perturbations), adversarial attacks (e.g., jailbreak attacks, prompt injection attacks), and out-of-distribution data. We also emphasize the importance of considering robustness in LLMs instruction—following ability, ensuring that models do not exhibit excessive safety, thereby compromising its performance on some normal tasks(e.g., "How to kill a process?"). In the realm of safety evaluation for LLMs, many researchers have proposed innovative methods and insights to evaluate robustness.
@@
Characterizing and Evaluating: Jailbreak attacks, particularly those involving carefully crafted prompts, pose significant challenges to the safety of LLMs. Chen et al. [53] introduces a comprehensive evaluation framework to mitigate this issue. Specifically, the researchers first construct a dataset covering 61 different harmful categories as the basis for evaluation. They then conduct a thorough evaluation of 10 advanced jailbreak attack methods against 13 popular LLMs. In evaluating the outputs of the LLMs, the authors consider a wide range of evaluation metrics, including Attack Success Rate (ASR), toxicity, fluency, grammatical errors, and token length, to facilitate an in-depth exploration of LLM safety when facing with jailbreak attacks. The results indicate that none of the evaluated LLMs exhibit initial resistance to harmful queries, with Vicuna and Mistral being the most vulnerable to jailbreak attacks within the LLMs family. Meanwhile, GPT-4 demonstrates the best performance across all categories, while Llama3 also shows strong capabilities. Among the jailbreak attack methods, ReNeLLM [134] achieves the highest average ASR, while ICA [132] and Cipher [133] exhibit lower performance. Furthermore, the correlation between evaluation metrics reveals that ASR is generally positively correlated with toxicity. Additionally, grammatical errors are positively correlated with token length, while fluency does not show a strong correlation with either grammatical errors or token length.
@@
DecodingTrust: [136] conducts a comprehensive trustworthiness evaluation of GPT-4 and GPT-3.5, focusing on adversarial robustness, out-of-distribution (OOD) robustness, and robustness against adversarial demonstration. To evaluate robustness on adversarial text attacks, the researchers construct three scenarios: (1) standard benchmark AdvGLUE [141], (2) AdvGLUE under various instructive task descriptions and system prompts, and (3) the more challenging adversarial text generation benchmark, AdvGLUE++. For OOD robustness, they explore scenarios including (1) inputs deviating from common training text styles, (2) recent events outside GPT models’ training data collection period, and (3) demonstrations with diverse OOD styles and domains added through in-context learning. For adversarial demonstration robustness, the evaluations involve injecting counterfactual examples, spurious correlations, and backdoors into the demonstrations to observe model’s performance. Overall, while GPT-4 typically outperforms GPT-3.5 on standard benchmarks, it proves more vulnerable to jailbreak attacks when specific system or user prompts are provided. The authors attribute this vulnerability to GPT-4’s tendency to follow misleading instructions more precisely.
@@
ASSERT: [12] proposes the Automatic Safety ScEnario Red Team (ASSERT) for robustness and safety evaluation of LLMs. ASSERT employs three novel methods—semantically aligned augmentation, targeted bootstrapping, and adversarial knowledge injection—to generate new test cases that explore robustness within language models. Semantically aligned augmentation aims to create samples with semantically equivalent but differently phrased expressions. Targeted bootstrapping generates new synthetic samples that are related to existing ones but not equivalent, while adversarial knowledge injection involves injecting adversarial knowledge during model inference. The researchers then evaluate LLMs on question-answering tasks across four critical AI application domains: outdoor, medical, household, and extra, requiring the models to determine whether a specific behavior $a$ should be implemented within the context $c$. The evaluation results reveal performance differences between semantically similar scenarios, with the model exhibiting instability up to a divergence of 11% in absolute classification accuracy. Adversarial attacks achieve error rates of 19.76% in zero-shot settings and 51.55% in adversarial four-shot demonstration settings, emphasizing the importance of few-shot demonstration.
JailbreakBench: [11] points out that current jailbreak evaluations lack clear practical standards, use incomparable computational costs and success rates for evaluation, and many works are difficult to reproduce. To address these challenges, the researchers introduce an open-source benchmark called JailbreakBench. First, JailbreakBench releases a jailbreak dataset of size 100, drawing from previous works across five aspects: behavior, goal (harmful queries), target (positive responses to harmful queries), category (based on OpenAI’s usage policy), and source (source dataset). Additionally, they maintain a repository of jailbreak artifacts and establish a standardized evaluation framework. This framework not only supports the definition of threat models, system prompts, and chat templates but also integrates a diverse range of evaluation methods, including rule-based string matching, GPT-4, and several models specifically designed for jailbreak evaluation (e.g., Llama Guard [137]). Evaluation of existing jailbreak attacks indicates that even recent closed-source, unprotected models are highly vulnerable to jailbreak attacks. Prompts with RS [138] outperform others such as PAIR [167], GCG [7], and Jailbreak Chat [139], achieving an attack success rate of up to 78% on GPT-4, highlighting the necessity for further defensive measures.
HarmBench: [10] designs a new red teaming attack and defense benchmark called HarmBench to address the lack of a standardized evaluation framework in automated red teaming. HarmBench consists of a set of harmful behaviors and an evaluation pipeline. These harmful behaviors include both textual and multimodal actions, and to enhance the robustness of the evaluation, HarmBench includes official validation/testing set partitions. Each behavior is categorized into semantic and functional categories. The semantic category describes the types of harmful behaviors (e.g., Cybercrime & Unauthorized Intrusion, Chemical and Biological Weapons/Drugs, etc.), including seven types. The functional category highlights unique properties of the behaviors that enable measuring the robustness of the target LLMs, including standard behaviors, copyright behaviors, contextual behaviors, and multimodal behaviors, totaling four types. In the evaluation pipeline, HarmBench ensures breadth by converting a series of behaviors into test cases. In addition, it employs fine-tuned Llama-2-13B and a hash-based classifier to evaluate the final results for non-copyright and copyright behaviors, respectively. The researchers compile 18 red teaming methods from 12 papers, and the experimental results indicate that no attack or defense is consistently effective, suggesting that robustness is independent of model size. This finding contradicts previous research [140], which points that larger models are more difficult to red team.

### v2

@@ II-C Importance
The importance of LLMs safety evaluation lies not only in ensuring reliable performance, but also in addressing the distinctive challenges that arise from the generative and open-ended nature of these models. Unlike traditional AI systems whose risks are often confined to narrow domains, LLMs operate across virtually unlimited contexts and user intentions, which amplifies both the scale and unpredictability of potential harms. A single unsafe response may propagate misinformation at scale, enable malicious use, or erode public trust, creating risks that extend far beyond technical failures [126, 127, 128].
The evaluation methods for LLMs’ general capabilities, such as accuracy benchmarks or instruction following tests, are not sufficient for safety evaluation, as safety concerns extend to subtle dimensions including refusal behavior, toxicity avoidance, bias mitigation, and resistance to adversarial prompting. Systematic safety evaluation is therefore indispensable for three reasons:
  • Managing Unique Risk Profiles. LLMs are open to the entire human society and face diverse and complex security risks. The risk of malicious use can occur in all possible scenarios. Evaluation enables the early detection of such vulnerabilities before real-world deployment.
  • Enabling Accountability and Governance. Transparent and standardized evaluation provides the evidence base necessary for legal compliance, policy development, and ethical oversight, particularly in high-stakes fields such as healthcare, law, and education.
  • Driving Technical Progress. Safety evaluation not only identifies shortcomings but also guides iterative model improvements, serving as a feedback loop for refining architectures, training strategies, and alignment techniques. This makes evaluation a scientific instrument for advancing both reliability and innovation.
In this survey, we emphasize that safety evaluation is not a peripheral safeguard but a central research agenda: it bridges societal expectations, regulatory demands, and technical innovation. By synthesizing safety tasks, datasets, benchmarks and methodologies, we aim to provide a roadmap that helps the community move beyond ad hoc testing toward standardized, rigorous, and impactful evaluation practices.
@@ III WHAT TO EVALUATE
This section provides an overview of various perspectives for evaluating the safety of LLMs, including toxicity, robustness, ethics, bias and fairness, truthfulness, and other specific tasks. It aims to illustrate what specific dimensions should be considered in evaluating and demonstrating the safety of LLMs, offering guidance for conducting comprehensive safety evaluations.
@@ III-A Toxicity
Toxicity refers to the presence of hate speech, offensive or abusive language, pornographic content, and other harmful expressions in text generated by LLMs. Such outputs not only pose risks of psychological harm to individual users but may also trigger broader social conflicts. As a result, evaluating the potential of LLMs to generate toxic content has become a critical step in ensuring their safe deployment. Existing efforts in this area can be broadly categorized into three lines of research: benchmark construction, adversarial and red-teaming based evaluation, and contextual or granular assessments.
III-A1 Benchmark Construction for Toxicity Evaluation
A major line of research has focused on building diverse and scalable benchmarks to capture toxicity in various settings. For example, ToxiGen [20] highlights that existing toxicity detectors often over-rely on surface-level mentions of minority identities, leading to the neglect of subtle hate speech and the over-detection of benign expressions. To address this issue, the authors developed a large-scale dataset generated via an adversarial classifier-in-the-loop decoding algorithm (ALICE), enabling the creation of implicit toxic texts across multiple demographic groups. ALICE adjusts the toxicity of the generated text by comparing a toxicity classifier with the text generator during beam search decoding. This adaptive control mechanism enables ToxiGen to not only generate texts that closely mimic human - written toxicity in nuanced expressions — including subtle biases toward underrepresented demographics — but also produce outputs that elude accurate classification by existing toxicity detectors, outperforming all prior human-curated corpora in both naturalistic mimicry and classifier evasion.
@@ III-A1 Benchmark Construction for Toxicity Evaluation
Similarly, ToxiChat [21] aims to address the challenge of detecting toxic content in real-world user-AI interactions. They note that most existing research on toxicity detection relies on benchmarks built from social media content, failing to capture the unique challenges of real-world conversations with AI systems. By collecting real user queries directed at open-source chatbots and introducing an uncertainty-guided human–AI annotation pipeline, the authors constructed a benchmark that reflects the safety concerns of interactive systems. The annotation pipeline uses the moderation API [91] to filter out low-confidence toxicity data, and the remaining data is manually reviewed, which greatly reduces the annotation workload. Furthermore, a systematic evaluation revealed that due to the domain inconsistency between the training data and real-world user prompts for chatbots, existing models not only fail to generalize to ToxiChat but also perform significantly worse on adversarial jailbreaking cases.
@@
Moving toward implicit harmful content, LifeTox [22] focuses on life-advice scenarios, where toxicity refers to a deeper, more concealed form of toxicity conveyed through linguistic features such as euphemisms, sarcasm, circumlocution, and metaphors, as well as extralinguistic knowledge like commonsense knowledge, world knowledge and social norms [129]. The authors argue that existing safety benchmark datasets are limited in diversity, primarily relying on red-team prompts, which often result in predictable and repetitive scenarios, failing to capture the complexity and variety of real-world situations. To address this gap, the dataset, constructed from Reddit forums, improves the detection of implicit harmful advice and demonstrates that fine-tuned RoBERTa[130] can rival LLMs in harmful content classification.
@@
Extending beyond English and Chinese, Jain et al. [23] propose PolygloToxicityPrompts, a large-scale multilingual benchmark covering 425K prompts across 17 languages. The prompts were generated by scraping documents from extensive corpora, with an average length of 400 GPT-4 tokens. Experimental results indicate significant differences in toxicity levels among LLMs across different languages, with an observed increase in toxicity as model size grows within the same family. Furthermore, models after instruction and preference-tuning exhibit lower toxicity compared to baseline models; However, the choice of preference-tuning method have a minimal impact on the models’ toxicity levels. The authors emphasize the need for further research on multilingual toxicity mitigation and the influence of model hyperparameters on toxicity.
III-A2 Adversarial and Red-Teaming Evaluation
@@
Another strand of work emphasizes adversarial attacks and red-teaming to probe the vulnerabilities of LLMs. Bhardwaj et al. [24] introduce RED-EVAL, a red-teaming benchmark that applies jailbreak attacks through a Chain of Utterances (CoU)-based prompt, facilitating a dialogue between two agents: a harmful agent Red-LM and a unsafe-helpful agent Base-LM. A harmful question is posed as the discourse for the Red-LM, and the model is instructed to generate a response according to the guidelines outlined in the prompt. They further propose RED-INSTRUCT, which enhances LLM safety through a two-phase approach: first, the CoU prompts are used to collect a dataset of harmful questions, termed HARMFULQA; Second, the model is safety-aligned by minimizing the negative log-likelihood of helpful responses while penalizing harmful ones. Experiments demonstrate that RED-EVAL achieves jailbreak success rates of 65% on GPT-4 and 73% on ChatGPT, underscoring the persistent risks of harmful content generation.
Similarly, Zhuo et al. [26] conducted a systematic red team evaluation of ChatGPT using the RealToxicPrompts dataset [131] and Perspective API [91]. The results showed that ChatGPT, due to its use of a cleaner training dataset, demonstrated superior performance compared to baseline models. However, the model remains vulnerable to prompt injection attacks, which can trigger toxic responses, and currently lacks mechanisms for early-stage toxicity detection.
III-A3 Contextual and Granular Toxicity Evaluation
Beyond benchmarks and adversarial probing, some works examine toxicity at different levels of granularity. FFT [25] investigate both utterance-level toxicity, where toxicity refers to literally-toxic language with some typical words, and context-level toxicity, where otherwise benign statements become harmful when interpreted in a broader conversational context. Their results show that LLMs generally exhibit higher toxicity rates at the contextual level, often by inadvertently affirming harmful issues despite generating literally non-toxic responses. This finding highlights the limitations of existing detectors in capturing nuanced, situational forms of toxicity.
Despite substantial progress, current approaches to toxicity evaluation face several unresolved challenges. First, the cultural and domain coverage of existing benchmarks remains limited, as most datasets are concentrated in a small set of languages or narrowly defined interaction scenarios, restricting their generalizability. Second, the emergence of increasingly sophisticated forms of implicit toxicity, conveyed through subtle linguistic cues or contextual framing, continues to outpace the ability of current models and detectors to capture them with sufficient accuracy. Third, the absence of standardized evaluation metrics and unified frameworks makes it difficult to compare results across studies, thereby hindering cumulative progress.
@@ III-B Robustness
Robustness refers to the ability of LLMs to maintain stable and safe outputs when exposed to noisy inputs, adversarial attacks such as jailbreak or prompt injection, and out-of-distribution (OOD) data. Beyond merely resisting malicious prompts, robustness also entails preserving instruction-following ability without exhibiting exaggerated safety, since models that indiscriminately refuse queries—regardless of their harmfulness (e.g., “How to kill a process?”)—sacrifice usability. Recent studies on robustness evaluation can be broadly grouped into three directions: adversarial and red-teaming evaluations, benchmarking and standardized frameworks, and the balance between instruction-following and excessive safety.
III-B1 Adversarial and Red-Teaming Evaluation
@@
Several works focus on systematically characterizing the vulnerabilities of LLMs against adversarial inputs. Jailbreak attacks, particularly those involving carefully crafted prompts, pose significant challenges to the safety of LLMs. Chen et al. [27] construct a dataset spanning 61 harmful categories to evaluate 10 jailbreak attack methods across 13 popular LLMs. Their framework employs multiple metrics, including attack success rate (ASR), toxicity, fluency, grammatical errors, and token length, enabling a multidimensional analysis of LLM safety. Results reveal that none of the evaluated models demonstrate inherent resistance to harmful queries, with Vicuna and Mistral among the most vulnerable, while GPT-4 and Llama3 exhibit comparatively stronger performance. Among the jailbreak attack methods, ReNeLLM [132] achieves the highest average ASR, while ICA [133] and Cipher [134] exhibit lower performance. They further observe that ASR correlates positively with toxicity. Additionally, grammatical errors are positively correlated with token length, while fluency does not show a strong correlation with either grammatical errors or token length.
@@
Complementing this work, DecodingTrust [28] conducts a comprehensive evaluation of GPT-4 and GPT-3.5, examining robustness to adversarial text attacks, OOD inputs, and adversarial demonstrations. For adversarial text attacks, the researchers construct three scenarios: (i) standard benchmark AdvGLUE [135], (ii) AdvGLUE under various instructive task descriptions and system prompts, and (iii) the more challenging adversarial text generation benchmark, AdvGLUE++. For OOD inputs, they explore scenarios including (i) inputs deviating from common training text styles, (ii) recent events outside GPT models’ training data collection period, and (iii) demonstrations with diverse OOD styles and domains added through in-context learning. For adversarial demonstrations, the evaluations involve injecting counterfactual examples, spurious correlations, and backdoors into the demonstrations to observe model’s performance. Overall, they find that GPT-4, though outperforming GPT-3.5 on standard benchmarks, is paradoxically more vulnerable to jailbreak attacks under certain prompts, owing to its stronger tendency to follow misleading instructions.
@@
Moving toward automated red-teaming, ASSERT [29] introduces three methods—semantically aligned augmentation, targeted bootstrapping, and adversarial knowledge injection—to generate new test cases. Semantically aligned augmentation aims to create samples with semantically equivalent but differently phrased expressions. Targeted bootstrapping generates new synthetic samples that are related to existing ones but not equivalent, while adversarial knowledge injection involves injecting adversarial knowledge during model inference. The researchers then evaluate LLMs on question-answering tasks across four critical AI application domains: outdoor, medical, household, and extra, requiring the models to determine whether a specific behavior $a$ should be implemented within the context $c$. Their results highlight that LLMs exhibit instability even under semantically similar scenarios, with divergences up to 11% in classification accuracy, while adversarial demonstrations induce error rates exceeding 50% in some settings. Collectively, these studies emphasize that LLM robustness remains fragile when confronted with diverse adversarial strategies and distributional shifts.
III-B2 Benchmarking and Standardized Frameworks
Recognizing the limitations of fragmented evaluations, another line of research develops systematic benchmarks and frameworks. JailbreakBench [30] addresses the lack of clear standards and reproducibility in jailbreak testing by two core means: first, releasing a benchmark dataset consisting of 100 prompts that cover behavior, harmful goal, positive responses, harmful category, and data source; second, maintaining a repository of jailbreak artifacts and establishing a standardized evaluation framework. Its standardized evaluation pipeline incorporates threat models, system prompts, chat templates and multiple evaluation methods, ranging from rule-based string matching to GPT-4-as-a-judge and several models specifically designed for jailbreak evaluation (e.g., Llama Guard [97]). Experiments indicate that even state-of-the-art models remain highly susceptible. Prompts with RS [136] outperform others such as PAIR [95], GCG [8], and Jailbreak Chat [137], achieving an ASR of up to 78% on GPT-4.
In parallel, HarmBench [31] establishes a more comprehensive red-teaming benchmark that categorizes harmful behaviors both semantically (e.g., cybercrime, bioweapons) and functionally (e.g., contextual, copyright, multimodal). In the evaluation pipeline, HarmBench ensures breadth by converting a series of behaviors into test cases. In addition, it employs fine-tuned Llama-2-13B and a hash-based classifier to evaluate the final results for non-copyright and copyright behaviors, respectively. Finally, it incorporates a total of 18 red-teaming methods from 12 studies, revealing that no attack or defense is universally effective, and that robustness does not correlate with model size—a finding that contradicts earlier assumptions [138]. Together, these benchmarks enhance comparability and reproducibility, while underscoring the persistent vulnerability of LLMs.
III-B3 Instruction-Following vs. Exaggerated Safety